eight more than three times a number is negative nineteen
the difference four times a number and nine is three
ten less than the product of two and a number is negative twelve

Found 2 solutions by oberobic, rfer:
Answer by oberobic(2304)

eight more than three times a number is negative nineteen:
.
3x+8 = -19
3x = -27
x = -9
.
the difference four times a number and nine is three
.
4x-9 = 3
4x = 12
x = 3
.
ten less than the product of two and a number is negative twelve
.
2x-10 = -12
2x = -2
x = -1
.
Done.

Answer by rfer(16322)

3x+8=-19
3x=-19-8
x=-27/3
x=-9
-----------
4x-9=3
4x=3+9
4x=12
x=12/4
x=3
-----------
2x-10=-12
2x=-12+10
2x=-2
x=-2/2
x=-1